OpenStudy (anonymous):

how will you simplify the radical of the square root 64

OpenStudy (anonymous):

the radical of the square root of 64 would be written like √√64 (radical and square root mean the same thing) the first radical would reduce √64 to 8, so it's now √8 but 8 is not a perfect square, so you need to find a factor of 8 which is a perfect square for example, 4. so you'd write it like √4*2 now you can simplify that to 2√2, which would be your answer

OpenStudy (anonymous):

im sorry its 20

OpenStudy (anonymous):

Then you asked the wrong question, perhaps you could be more precise with what you're trying to understand?

OpenStudy (anonymous):

how do you simplify the radical of the square root 20

OpenStudy (anonymous):

The radical of the square root of 20? You mean \[ \sqrt{\sqrt{20}} \] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Just do what was said above. Factor out the only perfect square in 20: \[ \sqrt{ \sqrt{4\cdot 5}} = \sqrt{2\sqrt{5}} \]

OpenStudy (anonymous):

kay so will that be the answer